• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name X. Zheng

    Log entry time 23:58:10 on March20,2005

    Entry number 142051

    keyword=Swing shift summary



    16:20- ROC6&ROC5 problem, see halog entry #142026

    17:25-18:45 Hall in controlled access for SRC crate reset, see #142032;

    18:50 Hall back in power permit;
    19:13 Beam back
    19:25-19:35 problem with Hall A laser
    20:24 Switched from Hall A to Hall C laser;
    21:24-21:36: running at 30uA due to Hall B beam optimization;

    Beam trip rate: ~15/hour (after switching to C laser)

    run summary:
    run# start stop #events /comments
    2827 15:20 NA 100K /end of run had CODA problem
    2828 18:46 18:46 / NA /dummy run to check CODA
    2829 18:53 18:54 / NA /dummy run to check CODA
    2830 19:14 19:32 / 50K /lost beam at the end due to laser problem
    2831 19:38 19:56 / 50K /laser problem again!
    2832 20:24 21:14 / 200K /good run
    2833 21:14 22:08 / 200K /good run
    2834 22:08 22:53 / 200K /good run
    2835 22:54 23:38 / 200K /good run
    2836 23:39 next shift...